# footnotetext: * Equal contribution, Corresponding author

### v1

@@ 4.2 Three-Stages Training of VIPGuard
It is crucial to first enhance the model’s ability to recognize and utilize facial attributes to enable identity protection for VIP users. It is deficient for naive MLLMs to understand human facial features. To this end, we first fine-tune the pre-trained MLLM on $\mathcal{D}_{FA}$ (described in 3). We integrate LoRA [23] modules into the pre-trained MLLM and perform supervised fine-tuning using an autoregressive loss, as defined in Eq. 3:
$\centering L(\bm{\theta})=-\sum_{i=1}^{N}\log\left[p_{\bm{\theta}}(x_{i}\mid x_{i-1},E_{V}(I))\right],\@add@centering$ (3)
where $\bm{\theta}$ denotes the parameters of both the inserted LoRA modules and the original MLLM, $N$ is the length of the output prompt, $x_{i}$ is the $i$-th token to be predicted, $E_{V}$ represents the image encoder of the MLLM, and $I$ is the input image. After training, the MLLM can understand and then recognize facial attributes.
@@
As illustrated in Figure 1, although the fake face appears realistic at first glance, subtle discrepancies remain in its detailed facial attributes. Therefore, we reformulate personalized deepfake detection as a fine-grained face recognition problem centered on the protected target, where forgeries are detected through the MLLM’s reasoning over global identity features and detailed facial attributes. In this stage, we further fine-tune the MLLM on $\mathcal{D}_{ID}^{{general}}$ for fine-grained facial recognition across face pairs—each consisting of an arbitrary reference face $I{{ref}}$ and an arbitrary query face $I_{{query}}$—thereby equipping it with a foundational ability to identity reason. $\mathcal{D}_{ID}^{\text{general}}$ contains a large number of positive (same-identity) and negative (including both different-identity and forgery) face pairs, each annotated with VQA-style reasoning questions and answers. These annotations support fine-grained identity discrimination by reasoning over both global and local facial prior. Here, we employ face recognition models [57, 13, 10] to provide face similarity scores, serving as a global facial prior, while the local facial prior—i.e., knowledge of facial attributes—has already been incorporated into the fine-tuned MLLM during Stage 1. Specifically, as shown in Figure  4, the reference image $I_{ref}$ and the query image $I_{query}$ are first inputted into the visual encoder $E_{V}$ of our MLLM to obtain the visual token sequences $f_{ref}$ and $f_{query}$, which can be formulated as follows
$\centering f_{ref}=E_{V}(I_{ref}),f_{query}=E_{V}(I_{query}).\@add@centering$ (4)
Pre-trained MLLMs typically lack task-specific optimization for facial recognition, rendering the vanilla visual features $f_{ref}$ and $f_{query}$ suboptimal for facial identity discrimination. Hence, we introduce a Cross-Attention (Cross-Attn) module [37] to capture fine-grained differences in visual tokens between distinct facial images, which is illustrated in Figure 4.
@@
$\centering L(\bm{{\theta}})=-\sum_{i=1}^{N}\log\left[p_{\bm{{\theta}}}(x_{i}\mid x_{i-1},E_{V}(I_{ref}),E_{V}(I_{query}),g)\right].\@add@centering$ (5)
@@
Upon completing Stage 2, the MLLM acquires a basic capability for fine-grained facial identity comparison with reasoning. In Stage 3, we further introduce a user-specific customization to enable personalized forgery detection for a given VIP user. To this end, we construct a dataset $\mathcal{D}_{ID}^{{vip}}$ with the same structure as $\mathcal{D}_{ID}^{{general}}$, centered on the VIP’s identity, allowing the model to identify suspicious images by reasoning based on the VIP user’s facial prior. Motivated by the Yo’LLaVA [42], to facilitate lightweight deployment, we incorporate a learnable VIP token trained on $\mathcal{D}_{ID}^{vip}$, denoted as $\bm{\mu}$, which encodes identity-specific features of the VIP user. During this stage, the parameters of the MLLM from Stage 2 are frozen, and only the VIP token is trained on $\mathcal{D}_{ID}^{{vip}}$, thereby refining the model’s ability to perceive and distinguish the target VIP user. As illustrated in Figure 4, the learnable VIP token $\bm{\mu}$, which is a vector of size $32\times d$, is employed to substitute the visual feature representation $f_{ref}$ of the reference image $I_{ref}$. Formally, the prediction procedure can be described as follows
$\centering L(\bm{\mu})=-\sum^{N}_{i}log[p_{\theta}(x_{i}|x_{i-1},\bm{\mu},E_{V}(I_{query}))].\@add@centering$ (6)
Notably, the reference images in $\mathcal{D}_{ID}^{{vip}}$ are not used; only the query image $I_{query}$ and the corresponding reasoning annotation are fed into the MLLM. After training, the MLLM can perform personalized forgery detection for the VIP user by determining whether the identity of query image $I_{{query}}$ belongs to the target individual and providing a user-centric explanation. Furthermore, during inference, we can substitute the corresponding VIP token $\bm{\mu}$ for each user, enabling efficient and accurate detection of malicious facial forgeries without requiring any model retraining. This design supports lightweight and scalable deployment.
@@ 5 Experiments
For evaluation, we report three standard metrics: Area Under the Curve (AUC), Equal Error Rate (EER), and Accuracy (ACC). AUC measures the model’s ability to distinguish between positive and negative classes across all thresholds. EER represents the point where the false acceptance rate is equal to the false rejection rate. ACC denotes the proportion of correct predictions. We use AUC and EER to compare our method following prior work [67, 70]. Moreover, ACC is used to compare with API-based commercial generators (e.g., GPT-4o [^5] GPT4o API version in use: GPT-4o-2024-08-06 [43]).
5.1 Evaluations on the proposed VIPEval and Other Existing Benchmarks
@@
We evaluated general deepfake detectors, ID-aware detectors, and our proposed method on VIPEval, reporting results for face swapping (Table  1) detection and entire faces synthesis detection(Table  2). Our method significantly improves the detection of all facial forgery types. General deepfake detectors (e.g., Effort [66]) excel at face swapping detection but fail to detect more realistic forgeries from commercial APIs due to unseen artifacts. ID-aware detectors, unlike general detectors reliant on low-level cues, leverage identity-related semantic consistency for robust detection across diverse generation techniques and data. However, existing ID-aware methods (e.g., DiffID, ICT-Ref) struggle with fully synthesized faces due to their reliance on global facial features, a limitation amplified by advancing forgery techniques. Conversely, VIP-Guard captures identity-specific cues by jointly leveraging global facial representations and local attribute analysis, enabling robust protection of VIP users with few real photos across diverse forgery methods.
@@
Methods BlendFace [52] Ghost3 [17] HifiFace [60] InSwap [27] MobileSwap [64] SimSwap [5] UniFace [62]
Xception [8] 53.89 / 46.59 61.08 / 42.27 71.70 / 34.09 64.79 / 38.64 98.12 / 7.50 64.91 / 38.41 59.91 / 42.50
EfficientNet [53] 33.94 / 60.91 43.52 / 54.32 61.22 / 42.95 37.34 / 58.64 89.67 / 64.45 56.55 / 45.00 46.53 / 50.23
UCF [69] 64.31 / 38.18 65.92 / 35.68 65.62 / 36.82 66.02 / 34.32 87.33 / 20.00 63.11 / 36.82 67.25 / 34.32
ProDet [7] 59.84 / 42.95 53.18 / 47.73 56.48 / 46.36 35.91 / 59.32 73.01 / 34.32 56.27 / 45.68 49.34 / 50.00
RECCE [3] 56.11 / 45.00 61.53 / 39.54 56.93 / 45.91 58.85 / 43.41 88.40 / 19.32 60.05 / 42.95 63.01 / 39.10
CDFA [39] 59.86 / 45.23 70.50 / 36.82 85.43 / 24.09 70.92 / 36.59 98.75 / 5.00 71.22 / 35.91 73.30 / 34.77
RepDFD [38] 70.34 / 35.00 78.76 / 28.18 80.09 / 27.50 71.66 / 34.32 95.76 / 10.45 79.40 / 27.95 82.36 / 25.45
Effort [66] 91.87 / 17.61 95.28 / 13.07 96.60 / 10.23 85.53 / 23.86 96.23 / 13.07 97.16 / 8.52 92.48 / 13.64
DiffID [73] 83.33 / 24.47 66.02 / 38.38 87.23 / 20.33 66.22 / 38.06 75.71 / 30.89 72.03 / 33.91 83.30 / 24.40
ICT-Ref [15] 88.67 / 13.37 86.52 / 15.51 85.90 / 14.05 84.34 / 17.87 87.45 / 13.52 86.57 / 14.65 82.73 / 19.04
@@
ConsistentID [25] Arc2Face [44] PuLID [22] GPT-4o [43] Jimeng AI [29] TongYi [55] Kling AI [31]
@@
UCF [69] 62.16 / 40.91 56.62 / 46.36 54.16 / 46.36 59.06 / 42.73 71.08 / 32.78 82.38 / 22.73 63.31 / 40.23
ProDet [7] 63.68 / 37.95 59.62 / 40.91 67.02 / 36.82 59.23 / 40.91 59.71 / 42.73 89.80 / 18.41 72.53 / 32.73
RECCE [3] 68.56 / 36.82 57.79 / 47.72 63.85 / 40.00 83.63 / 22.73 69.00 / 35.23 97.00 / 7.50 70.94 / 37.72
CDFA [39] 77.62 / 30.00 67.09 / 39.09 67.93 / 37.50 73.46 / 32.73 71.98 / 34.09 90.32 / 17.95 77.47 / 30.00
RepDFD [38] 83.52 / 24.32 61.67 / 41.59 74.65 / 32.27 73.62 / 32.27 62.78 / 40.91 93.80 / 14.09 60.10 / 43.41
Effort [66] 58.68 / 47.35 57.03 / 44.89 56.31 / 44.89 49.63 / 48.30 63.60 / 40.34 82.93 / 23.86 56.84 / 44.89
DiffID [73] 75.85 / 32.01 78.47 / 28.86 70.36 / 35.72 45.26 / 52.97 64.29 / 39.35 84.66 / 23.58 69.51 / 35.56
ICT-Ref [15] 63.15 / 39.84 70.27 / 32.84 72.36 / 31.93 58.59 / 40.93 65.36 / 35.73 74.88 / 24.41 50.05 / 45.98
@@
GPT-4o-2024-08-06[43] 71.85 84.36 94.35 70.72 46.70 50.90
Gemini-2.5-pro-exp-03-25 [20] 80.87 82.68 91.75 83.58 69.85 79.98
Qwen2.5VL 7B [1] 50.05 49.86 49.99 50.02 49.51 49.69
LLaMA3.2Vision 11B [21] 44.76 43.93 45.71 54.09 38.56 43.61
FFAA [26] 88.60 89.95 91.36 89.30 63.53 59.89
@@
Comparison with Other LLM-based Methods on VIPEval
This experiment evaluated the detection capabilities of Multimodal Large Language Models (MLLMs) on VIPEval by comparing various MLLM-based methods, including FFAA [26], an MLLM specialized for face forgery detection, alongside naive MLLMs. Detection performance was measured using Accuracy (ACC) due to the API’s binary (Real/Fake) output. As presented in Table  3, our method consistently outperforms other approaches across all forged image types, demonstrating the effectiveness of identity-specific semantic detection.
Table 4: Evaluation of frame-level performance (AUC (%) / EER (%)) in CelebDF [36] and DF40 [67] under one-shot setting. For each identity, only a single real image is available.
@@
Figure 5: Visual illustration of the analysis of VIP-Guard detecting anomalous local facial attributes for EFS (left) and FS (right).
One-shot Performance on other Deepfake detection datasets
This experiment evaluates our method against several existing ID-aware approaches on established benchmark datasets. The evaluation specifically employed the challenging CelebDF [36] dataset and a CelebDF-related subset of DF40 [67]. Owing to limited real sample diversity in these datasets, our method was evaluated in a one-shot setting, using a single real image per user. To mitigate similarity from shared video sources, reference and test images were sampled from different videos. As shown in Table 4, our method still achieves competitive performance compared to other approaches. Notably, due to limited real image availability, these results were obtained using VIP-Guard at Stage 2, omitting Stage 3. The results also verify the effectiveness of Stage 2 of our method.

### v2

@@ 4.3 Three-Stages Training of VIPGuard
It is crucial first to enhance the model’s capability to recognize and utilize fine-grained facial attributes, as naive MLLMs inherently lack a sufficient understanding of human facial features for effective VIP identity protection. To this end, we fine-tune the pre-trained MLLM on $\mathcal{D}_{FA}$, which contains a large number of samples for facial attribute recognition (see Section 3). We integrate LoRA [lora] modules into the pre-trained MLLM and perform supervised fine-tuning using an autoregressive loss, as defined in Eq. 3:
$\centering L(\bm{\theta})=-\sum_{i=1}^{N}\log\left[p_{\bm{\theta}}(x_{i}\mid x_{<i},E_{V}(I))\right],\@add@centering$ (3)
where $\bm{\theta}$ denotes the parameters of both the inserted LoRA modules and the original MLLM, $N$ is the length of the output prompt, $x_{i}$ is the $i$-th token to be predicted while $x_{<i}$ are the previous tokens, $E_{V}$ represents the visual encoder of the MLLM, and $I$ is the input image. After training, the MLLM can understand and then recognize facial attributes.
@@
As illustrated in Figure 1, although the fake face appears realistic at first glance, subtle discrepancies remain in its detailed facial attributes. Therefore, we reformulate personalized deepfake detection as a fine-grained face recognition problem centered on the protected target, where forgeries are detected through the MLLM’s reasoning over global identity features and detailed facial attributes. In this stage, we further fine-tune the MLLM on $\mathcal{D}_{ID}^{{general}}$ for fine-grained facial recognition across face pairs—each consisting of an arbitrary query face $I_{{query}}$ and an arbitrary input face $I_{{Img}}$—thereby equipping it with a foundational ability to identity reason. $\mathcal{D}_{ID}^{\text{general}}$ contains a large number of positive (same-identity) and negative (including both different-identity and forgery) face pairs, each annotated with VQA-style reasoning questions and answers. These annotations support fine-grained identity discrimination by reasoning over both global and local facial prior. Here, we employ face recognition models [cosface, arcface, transface] to provide face similarity scores, serving as a global facial prior, while the local facial prior—i.e., knowledge of facial attributes—has already been incorporated into the fine-tuned MLLM during Stage 1. Specifically, as shown in Figure  4, the query $I_{query}$ and the input image $I_{Img}$ are first inputted into the visual encoder $E_{V}$ of our MLLM to obtain the visual token sequences $f_{query}$ and $f_{Img}$. Pre-trained MLLMs typically lack task-specific optimization for facial recognition, rendering the vanilla visual features $f_{query}$ and $f_{Img}$ suboptimal for facial identity discrimination. Hence, we introduce a Cross-Attention (Cross-Attn) module [crossattn] to capture fine-grained differences in visual tokens between distinct facial images, which is formulated as follows
$\centering g=\mathrm{softmax}\!\left(\frac{QK^{\top}}{\sqrt{d_{K}}}\right)V,\quad\text{where }Q=f_{\text{query}},\;K=V=f_{\text{Img}}.\@add@centering$ (4)
@@
$\centering L(\bm{{\theta}})=-\sum_{i=1}^{N}\log\left[p_{\bm{{\theta}}}(x_{i}\mid x_{<i},f_{query},f_{Img},g)\right].\@add@centering$ (5)
@@
Upon completing Stage 2, the MLLM acquires a basic capability for fine-grained facial identity comparison with reasoning. In Stage 3, we further introduce a user-specific customization to enable personalized forgery detection for a given VIP user. To this end, we construct a dataset $\mathcal{D}_{ID}^{{vip}}$ with the same structure as $\mathcal{D}_{ID}^{{general}}$, centered on the VIP’s identity, allowing the model to identify suspicious images by reasoning based on the VIP user’s facial prior. Motivated by the Yo’LLaVA [nguyen2024yo], to facilitate lightweight deployment, we incorporate a learnable VIP token trained on $\mathcal{D}_{ID}^{vip}$, denoted as $\bm{\mu}$, which encodes identity-specific features of the VIP user. During this stage, the parameters of the MLLM from Stage 2 are frozen, and only the VIP token is trained on $\mathcal{D}_{ID}^{{vip}}$, thereby refining the model’s ability to perceive and distinguish the target VIP user. As illustrated in Figure 4, the learnable VIP token $\bm{\mu}$, which is a vector of size $32\times d$, is employed to substitute the visual feature representation $f_{query}$ of the query image $I_{query}$. Formally, the prediction procedure can be described as follows
$\centering g=\mathrm{softmax}\!\left(\frac{QK^{\top}}{\sqrt{d_{K}}}\right)V,\quad\text{where }Q=\bm{\mu},\;K=V=f_{\text{Img}}.\@add@centering$ (6)
$\centering L(\bm{\mu})=-\sum^{N}_{i=1}log[p_{\theta}(x_{i}|x_{<i},\bm{\mu},f_{query},g)].\@add@centering$ (7)
Notably, the query images in $\mathcal{D}_{ID}^{{vip}}$ are not used; only the input image $I_{Img}$ and the corresponding reasoning annotation are fed into the MLLM. After training, the MLLM can perform personalized forgery detection for the VIP user by determining whether the identity of the input image $I_{{Img}}$ belongs to the target individual and providing a user-centric explanation. Furthermore, during inference, we can substitute the corresponding VIP token $\bm{\mu}$ for each user, enabling efficient and accurate detection of malicious facial forgeries without requiring any model retraining. This design supports lightweight and scalable deployment.
@@ 5 Experiments
For evaluation, we report three standard metrics: Area Under the Curve (AUC), Equal Error Rate (EER), and Accuracy (ACC). AUC measures the model’s ability to distinguish between positive and negative classes across all thresholds. EER represents the point where the false acceptance rate is equal to the false rejection rate. ACC denotes the proportion of correct predictions. We use AUC and EER to compare our method following prior work [yan2024df40, DeepfakeBench_YAN_NEURIPS2023]. Moreover, ACC is used to compare with API-based commercial generators (e.g., GPT-4o [^4] GPT4o API version in use: GPT-4o-2024-08-06 [gpt4o]).
@@
We evaluated general deepfake detectors, ID-aware detectors, and our proposed method on VIPEval, reporting results for face swapping (Table  1) detection and entire faces synthesis detection (Table  2). Our method significantly improves the detection of all facial forgery types. General deepfake detectors (e.g., Effort [yan2024effort]) excel at face swapping detection but fail to detect more realistic forgeries from commercial APIs due to unseen artifacts. ID-aware detectors, unlike general detectors reliant on low-level cues, leverage identity-related semantic consistency for robust detection across diverse generation techniques and data. However, existing ID-aware methods (e.g., DiffID, ICT-Ref) struggle with fully synthesized faces due to their reliance on global facial features, a limitation amplified by advancing forgery techniques. Conversely, VIP-Guard captures identity-specific cues by jointly leveraging global facial representations and local attribute analysis, enabling robust protection of VIP users with few real photos across diverse forgery methods.
@@
Methods BlendFace [blendface] Ghost3 [ghost3] HifiFace [wang2021hififace] InSwap [inswapper] MobileSwap [xu2022mobilefaceswap] SimSwap [simswap] UniFace [uniface]
Xception [chollet2017xception] 53.89 / 46.59 61.08 / 42.27 71.70 / 34.09 64.79 / 38.64 98.12 / 7.50 64.91 / 38.41 59.91 / 42.50
EfficientNet [tan2019efficientnet] 33.94 / 60.91 43.52 / 54.32 61.22 / 42.95 37.34 / 58.64 89.67 / 64.45 56.55 / 45.00 46.53 / 50.23
UCF [yan2023ucf] 64.31 / 38.18 65.92 / 35.68 65.62 / 36.82 66.02 / 34.32 87.33 / 20.00 63.11 / 36.82 67.25 / 34.32
ProDet [cheng2024can] 59.84 / 42.95 53.18 / 47.73 56.48 / 46.36 35.91 / 59.32 73.01 / 34.32 56.27 / 45.68 49.34 / 50.00
RECCE [cao2022end] 56.11 / 45.00 61.53 / 39.54 56.93 / 45.91 58.85 / 43.41 88.40 / 19.32 60.05 / 42.95 63.01 / 39.10
CDFA [lin2024fake] 59.86 / 45.23 70.50 / 36.82 85.43 / 24.09 70.92 / 36.59 98.75 / 5.00 71.22 / 35.91 73.30 / 34.77
RepDFD [lin2025standing] 70.34 / 35.00 78.76 / 28.18 80.09 / 27.50 71.66 / 34.32 95.76 / 10.45 79.40 / 27.95 82.36 / 25.45
Effort [yan2024effort] 91.87 / 17.61 95.28 / 13.07 96.60 / 10.23 85.53 / 23.86 96.23 / 13.07 97.16 / 8.52 92.48 / 13.64
DiffID [diff_id] 83.33 / 24.47 66.02 / 38.38 87.23 / 20.33 66.22 / 38.06 75.71 / 30.89 72.03 / 33.91 83.30 / 24.40
ICT-Ref [ict_ref] 88.67 / 13.37 86.52 / 15.51 85.90 / 14.05 84.34 / 17.87 87.45 / 13.52 86.57 / 14.65 82.73 / 19.04
@@
ConsistentID [huang2024consistentid] Arc2Face [arc2face] PuLID [guo2024pulid] GPT-4o [gpt4o] Jimeng AI [jimengai] TongYi [TongYiWanXiang] Kling AI [kling]
@@
UCF [yan2023ucf] 62.16 / 40.91 56.62 / 46.36 54.16 / 46.36 59.06 / 42.73 71.08 / 32.78 82.38 / 22.73 63.31 / 40.23
ProDet [cheng2024can] 63.68 / 37.95 59.62 / 40.91 67.02 / 36.82 59.23 / 40.91 59.71 / 42.73 89.80 / 18.41 72.53 / 32.73
RECCE [cao2022end] 68.56 / 36.82 57.79 / 47.72 63.85 / 40.00 83.63 / 22.73 69.00 / 35.23 97.00 / 7.50 70.94 / 37.72
CDFA [lin2024fake] 77.62 / 30.00 67.09 / 39.09 67.93 / 37.50 73.46 / 32.73 71.98 / 34.09 90.32 / 17.95 77.47 / 30.00
RepDFD [lin2025standing] 83.52 / 24.32 61.67 / 41.59 74.65 / 32.27 73.62 / 32.27 62.78 / 40.91 93.80 / 14.09 60.10 / 43.41
Effort [yan2024effort] 58.68 / 47.35 57.03 / 44.89 56.31 / 44.89 49.63 / 48.30 63.60 / 40.34 82.93 / 23.86 56.84 / 44.89
DiffID [diff_id] 75.85 / 32.01 78.47 / 28.86 70.36 / 35.72 45.26 / 52.97 64.29 / 39.35 84.66 / 23.58 69.51 / 35.56
ICT-Ref [ict_ref] 63.15 / 39.84 70.27 / 32.84 72.36 / 31.93 58.59 / 40.93 65.36 / 35.73 74.88 / 24.41 50.05 / 45.98
@@
Comparison with Other LLM-based Methods on VIPEval
This experiment evaluated the detection capabilities of Multimodal Large Language Models (MLLMs) on VIPEval by comparing various MLLM-based methods, including FFAA [huang2024ffaa], an MLLM specialized for face forgery detection, alongside naive MLLMs. Detection performance was measured using Accuracy (ACC) due to the API’s binary (Real/Fake) output. As presented in Table  3, our method consistently outperforms other approaches across all forged image types, demonstrating the effectiveness of identity-specific semantic detection.
@@
GPT-4o-2024-08-06[gpt4o] 71.85 84.36 94.35 70.72 46.70 50.90
Gemini-2.5-pro-exp-03-25 [Gemini] 80.87 82.68 91.75 83.58 69.85 79.98
Qwen2.5VL 7B [bai2025qwen2] 50.05 49.86 49.99 50.02 49.51 49.69
LLaMA3.2Vision 11B [grattafiori2024llama] 44.76 43.93 45.71 54.09 38.56 43.61
FFAA [huang2024ffaa] 88.60 89.95 91.36 89.30 63.53 59.89
@@
One-shot Performance on other Deepfake detection datasets
This experiment evaluates our method against several existing ID-aware approaches on established benchmark datasets. The evaluation specifically employed the challenging CelebDF [li2019celeb] dataset and a CelebDF-related subset of DF40 [yan2024df40]. Owing to limited real sample diversity in these datasets, our method was evaluated in a one-shot setting, using a single real image per user. To mitigate similarity from shared video sources, reference and test images were sampled from different videos. As shown in Table 4, our method still achieves competitive performance compared to other approaches. Notably, due to limited real image availability, these results were obtained using VIP-Guard at Stage 2, omitting Stage 3. The results also verify the effectiveness of Stage 2 of our method.
Table 4: Evaluation of frame-level performance (AUC (%) / EER (%)) in CelebDF [li2019celeb] and DF40 [yan2024df40] under one-shot setting. For each identity, only a single real image is available.
@@
Figure 5: Visual illustration of the analysis of VIP-Guard detecting anomalous local facial attributes for EFS (left) and FS (right). The two real images are sourced from LAION-Face [laion_face], while the fake images in the left and right subfigures were generated by GPT-4o [gpt4o] and HifiFace [wang2021hififace], respectively.
5.1 Effectiveness in Annotation-Free Scenarios
To validate VIPGuard’s practicality, we tested its Stage 3 performance when trained using only VIP images, foregoing textual annotations. As shown in Table 5, the model’s performance remains strong, with the average AUC decreasing only slightly from 98.98% to 95.86%. This robustness stems directly from the design of our framework. Stage 2 is responsible for internalizing a vast repository of discriminative knowledge by learning to distinguish between arbitrary image pairs with the aid of textual supervision. This foundational pre-training is so effective that the User-specific fine-tuning in Stage 3 requires only visual information to achieve a high degree of accuracy. Consequently, this experiment validates VIPGuard as a highly practical and readily deployable framework for real-world applications, including annotation-free scenarios. The model’s ability to achieve such high efficacy under these constraints is a direct testament to the comprehensive and foundational knowledge provided by Stage 2.
Table 5: Performance (AUC (%)) of VIPGuard under different training configurations in Stage 3. Images + Annotation denotes training with both images and textual descriptions, while Only Images uses visual inputs only.
   Variant    BlendFace    InSwap    Arc2Face    PuLID    Average
   Only Images    98.45    92.91    94.35    97.72    95.86
   Images + Annotation    99.48    99.43    98.05    98.96    98.98